OpenStudy (anonymous):

3^2x-3=5^x+2 solve this equation in terms of the natural logarithm. Give an approximation to the nearest hundreth of a unit.

OpenStudy (anonymous):

\[ 3^{2x}-3=5^x+2 \]is this how it is?

OpenStudy (anonymous):

no the exponent is 2x-3 and x+2

OpenStudy (p0sitr0n):

add ln on both sides

OpenStudy (anonymous):

so, \[\Large{ 3^{2x-3}=5^{x+2}\\ 3^{2x}3^{-3}=5^x5^2\qquad\text{split the exponent}\\ \text{now, bring the 'x' terms together}\\ \frac{3^{2x}}{5^x}={5^2\over 3^{-3}}\\ \left(3^2\over 5\right)^x=5^23^3\\ \left(9\over 5\right)^x=25\times 27\ }\]now, take the logarithm

OpenStudy (anonymous):

when you take the logrithm now, the exponent "x" will come down by property

OpenStudy (anonymous):

supposed to be the natural logarithm

OpenStudy (anonymous):

you are free to use any logarithm base.

OpenStudy (anonymous):

oh ok well for this problem it asks to use natural

OpenStudy (anonymous):

then use it. there is another way for this too.. \[\Large{ 3^{2x-3}=5^{x+2}\\ \ln(3^{2x-3})=\ln(5^{x+2})\\ (2x-3)\ln(3)=(x+2)\ln(5) }\]then simplify this further

OpenStudy (anonymous):

I used the property \[\log(a^m)=m\times \log(a)\]

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

ya I dont know how to simplify it further though

OpenStudy (anonymous):

Im terrible with this stuff

OpenStudy (anonymous):

use a calculator to find the logs. on your calculator there should be a button that says "ln"=natural log and "log"=log base 10

OpenStudy (anonymous):

so take the ln of 3 and the ln of 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

then show me what you got

OpenStudy (anonymous):

(2x-3)(1.10)=(x+2)(1.61)

OpenStudy (anonymous):

now, we expand the paranthesis..

OpenStudy (anonymous):



OpenStudy (anonymous):

2.2x-3.3=1.61+3.22

OpenStudy (anonymous):

1.61x

OpenStudy (anonymous):

great.